# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v1

@@ IV The Proposed HFLDD
In this section, we will firstly introduce the key idea of the proposed framework: HFLDD. Then, we will introduce the details of each part and finally provide an analysis of the communication cost of HFLDD.
@@ IV-A Overview
The key idea of HFLDD is to construct heterogeneous client clusters, where the data labels among different clusters are balanced and approximate to the global distribution. The server can communicate with each cluster to obtain a global model similar to one achieved under IID data conditions.
The overall pipeline of our proposed HFLDD is shown in Fig. 1, and Table I summarizes the important notations. Briefly, HFLDD can be mainly divided into four parts: label knowledge collection, heterogeneous client clustering, IID dataset generation, and model training. Each part includes one or several steps. We will briefly introduce each part and the details will be shown in the following subsection.
  • Label knowledge collection: Based on pre-trained models, each client predicts soft labels of the data-points in a global dataset, then uploads all the soft labels to the server (Step 1). The server obtains the label knowledge from all the clients.
  • Heterogeneous client clustering: The server calculates the similarities among clients’ labels (Step 2) and groups clients into homogeneous clusters. Subsequently, cluster sampling is performed by the server to identify members of heterogeneous clusters and randomly selects one cluster header for each cluster. The clustering results are then distributed to the clients (Step 3).
  • IID dataset generation: All clients within a cluster, except the header, perform dataset distillation and send their distilled data to the cluster header (Step 4). As a result, The cluster header obtains a dataset that is approximately IID.
  • Model training: Each cluster header conducts local model training using a combined dataset, and uploads the local model parameters to the server. The server updates the global model and distributes it to cluster headers (Step 5). This step repeats until model convergence.
@@ IV-B Details
The details about the implementation of each part in HFLDD are introduced in this subsection.
@@ IV-B1 Label knowledge collection
Each client first trains a local model on their own dataset $\mathcal{D}_{i}$. Since the trained local model is driven by the client dataset, it learns relevant knowledge about the local data. Subsequently, we introduce a finite and globally available dataset $\mathcal{D}_{g}=\{\left(X_{g},y_{g}\right)\}$, where $X_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times d}$, $y_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}}$ and $|\mathcal{D}_{g}|$ is the number of samples in $\mathcal{D}_{g}$. The global dataset has a distribution different from the clients’ datasets and is used to generate soft labels for each client. Specifically, each client uses a globally available dataset as input for its locally pre-trained model to generate a two-dimensional tensor $S_{i}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}},\forall i=1,\cdots,N$. This tensor represents the soft labels of the global dataset based on client $i$’s local model. ${S}_{i,k}^{c}$ denotes the probability of sample $k$ belonging to class $c$. It has been verified that the above soft labels can reflect the label knowledge of the local dataset [30]. Therefore, after receiving the soft labels from the clients, the server obtains the label knowledge of each client without knowing the original dataset, which prevents data privacy leakage.
@@ IV-B2 Heterogeneous Client Clustering
Based on the similarity matrix, the server employs the K-Means algorithm to cluster all the clients, by treating each row of the similarity matrix as a data point. This process groups clients with similar label knowledge into homogeneous clusters $\mathcal{H}_{o}=\left\{\mathcal{H}_{o,1},\ldots,\mathcal{H}_{o,K}\right\}$. However, our goal is to construct heterogeneous clusters $\mathcal{H}_{e}=\left\{\mathcal{H}_{e,1},\ldots,\mathcal{H}_{e,H}\right\}$ with approximately IID intra-cluster data. Therefore, to construct heterogeneous clusters, we randomly select one client from each of the homogeneous clusters, and form a new cluster with them. The details are shown in Algorithm 1. Through this sampling process, we ensure that the newly formed clusters contain clients with different label knowledge. Fig. 2 illustrates the overall process of heterogeneous client clustering. Each client’s local data contains only one label, which makes the data among different clients are severe Non-IID. After clustering, each homogeneous cluster contains only one label, while each heterogeneous cluster contains labels from all the clients.
Then, the server randomly selects one client within each heterogeneous cluster as the header for this cluster, and pushes the cluster information to each client. In this way, the clients constructs a two-layer topology, as shown in Fig. 1
@@ IV-B3 IID dataset generation
In each heterogeneous cluster, the cluster header acts as an agent for all the cluster members, conducts local model training, and communicates with the server. Each cluster member, except the header, sends a distilled data of much smaller size to the cluster header. In this work, the clients employ the KIP method to obtain distilled data, as shown in Section III-A. For the $h$-th heterogenous cluster, when the cluster header aggregates the distilled data from the entire cluster, it obtains a dataset $\mathcal{D}_{h}$, whose distribution approximates the global distribution. We transform the problem into a learning process between the server and a number of cluster headers with approximately IID data.
@@ IV-B4 Model training
As we mentioned above, after transforming our problem into distributed learning with IID data, we can apply classical FL method, such as FedAvg[9], for model training. In communication round $t$, the server broadcasts the global model parameters $\omega^{t}$ to all the connected cluster headers. Subsequently, cluster header $h$ updates its local parameters to $\omega_{h}^{t}\,\,=\,\,\omega^{t}$, and performs $E_{2}$ epochs of local updates on $\mathcal{D}_{h}$. At the $(k+1)$-th epoch of local update, the local model parameters of cluster header $h$ is updated as
@@
where $h=1,\cdots,H$, $k=0,\cdots,E_{2}-1$, $\xi_{h}^{t,k}$ denotes a randomly selected batch from $\mathcal{D}_{h}$, $\eta$ denotes the learning rate and $\nabla F_{h}\left(\cdot\right)$ denotes the gradient of local loss function. Finally, the server aggregates the updated local models $\omega_{1}^{t,E_{2}},...,\omega_{H}^{t,E_{2}}$ from all the cluster headers to generate a new global model.
The overall process of the proposed HFLDD is illustrated in Algorithm 2.
IV-C Communication Cost Analysis
Now, we analyze the total communication cost of HFLDD. Considering that the initialization of the server model can be sent as a random seed [27], we can ignore the communication cost of the server sending random seeds. Thus, we only consider the resources consumed for uploading soft labels to the server at the stage of label knowledge collection, transmitting the distilled data within each cluster at the stage of IID dataset generation, and communication between the server and cluster headers at the stage of model training.
Let $a_{h}$ denote the cluster header of heterogeneous cluster $h$, and $\left|\tilde{\mathcal{D}_{i}}\right|$ denotes the size of the distilled dataset on client $i$. Suppose transmitting each parameter incurs a communication overhead of $B_{1}$ (usually 32 bits), and transmitting each distilled data incurs a communication overhead of $B_{2}$ (1 × 8 bits or 3 × 8 bits for each gray-scale or RGB pixel). For $N$ clients, the communication cost for uploading the soft label is $N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}$. The total communication cost for transmitting the distilled data within each cluster is $\sum_{h=1}^{H}{\sum_{\begin{subarray}{c}i\in\mathcal{H}_{e,h}\\ i\neq a_{h}\end{subarray}}{\left|\tilde{\mathcal{D}_{i}}\right|}}\cdot B_{2}$. At the model training stage, the total cost for the model communication between the server and cluster headers at $T$ rounds is $H\cdot|\omega|\cdot(2T-1)\cdot B_{1}$. Consequently, the total communication cost of HFLDD is
$Cost_{hfldd}$ $=N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}$
$~{}~{}~{}+H\cdot|\omega|\cdot(2T-1)\cdot B_{1}.$ (6)
@@ IV-C Communication Cost Analysis
19:     select cluster header $a_{h}$ randomly from $\mathcal{H}_{e,h}$
@@
V Experimental Evaluation

### v2

@@ IV The Proposed HFLDD
In this section, we will first introduce the key idea of the proposed framework HFLDD. Then, we will introduce the details of each part and finally provide an analysis of the communication cost of HFLDD.
@@ IV-A Overview
The key idea of HFLDD is to construct heterogeneous client clusters, where the data labels among different clusters are balanced and approximate the global distribution. The server can communicate with each cluster to obtain a global model similar to that achieved under IID data conditions.
The overall pipeline of the proposed HFLDD framework is illustrated in Fig. 1, which consists of four key stages: 1) label knowledge collection, 2) heterogeneous client clustering, 3) IID dataset generation, and 4) model training. Each stage is associated with a specific set of operations, as described below:
  1. 1.
    Label Knowledge Collection (Step 1): Each client independently trains a local model using its private non-IID dataset. These locally trained models are then used to infer soft labels on a globally available auxiliary dataset. The resulting soft-label matrices are uploaded to the server. This stage allows the server to gather label distribution information without accessing the raw private data, thus preserving client privacy.
  2. 2.
    Heterogeneous Client Clustering (Steps 2–3): The server computes pairwise similarities between clients based on their soft labels and groups the clients with similar label distributions into homogeneous clusters. To enhance diversity, the server then constructs heterogeneous clusters by randomly sampling one client from each homogeneous cluster, and selects one client from each heterogeneous cluster as the cluster head. The clustering results are distributed to all clients.
  3. 3.
    IID Dataset Generation (Step 4): Within each heterogeneous cluster, all clients except the cluster head perform dataset distillation to generate a compact synthetic dataset that approximates their local data distribution. These distilled datasets are sent to the cluster head. The head aggregates all received distilled data, including its own data, to form a more balanced, approximately IID dataset. This step helps mitigate statistical heterogeneity before training begins.
  4. 4.
    Model Training (Step 5): The server distributes the current global model to all cluster heads. Each head updates the model using the locally aggregated IID-like dataset and sends the updated model back to the server. The server aggregates the updated models using standard FL techniques such as FedAvg. This process is repeated for multiple rounds until model convergence.
@@ IV-B Details
The details of the implementation of each part in HFLDD are introduced in this subsection. Table I summarizes the important notations.
@@ IV-B1 Label knowledge collection
Each client first trains a local model on their own dataset $\mathcal{D}_{i}$. Since the trained local model is driven by the client dataset, it learns relevant knowledge about the local data. Subsequently, we introduce a finite and globally available dataset $\mathcal{D}_{g}=\{\left(X_{g},y_{g}\right)\}$  [^1] The global dataset $\mathcal{D}_{g}$ used for soft label generation can be a small, publicly available dataset related to the same task but it is not necessarily representative of the client data, where $X_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times d}$, $y_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}}$ and $|\mathcal{D}_{g}|$ is the number of samples in $\mathcal{D}_{g}$. The global dataset has a distribution different from the datasets of the clients and is used to generate soft labels for each client. Specifically, each client uses a globally available dataset as input for its locally pre-trained model to generate a two-dimensional tensor $S_{i}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}},\forall i=1,\cdots,N$. This tensor represents the soft labels of the global dataset based on the local model of the client $i$. ${S}_{i,k}^{c}$ denotes the probability that the sample $k$ belongs to class $c$. It has been verified that the above soft labels can reflect the knowledge of the label of the local dataset [32, 48]. Therefore, after receiving the soft labels from the clients, the server obtains the label knowledge of each client without knowing the original dataset, which prevents data privacy leakage.
@@ IV-B2 Heterogeneous Client Clustering
Based on the similarity matrix, the server employs the K-Means algorithm to cluster all the clients, by treating each row of the similarity matrix as a data point. This process groups clients with similar knowledge of labels into homogeneous groups $\mathcal{H}_{o}=\left\{\mathcal{H}_{o,1},\ldots,\mathcal{H}_{o,K}\right\}$. However, our goal is to construct heterogeneous clusters $\mathcal{H}_{e}=\left\{\mathcal{H}_{e,1},\ldots,\mathcal{H}_{e,H}\right\}$ with approximately intracluster IID data. Therefore, to construct heterogeneous clusters, we randomly select one client from each of the homogeneous clusters and form a new cluster with them. The details are shown in Algorithm 1. Through this sampling process, we ensure that the newly formed clusters contain clients with different label knowledge. Fig. 2 illustrates the overall process of heterogeneous client clustering. Each client’s local data contains only one label, which makes the data among different clients severe non-IID. After clustering, each homogeneous cluster contains only one label, while each heterogeneous cluster contains labels from all clients.
Then, the server randomly selects one client within each heterogeneous cluster as the head for this cluster, and pushes the cluster information to each client. In this way, the clients construct a two-layer topology, as shown in Fig. 1
@@ IV-B3 IID dataset generation
In each heterogeneous cluster, the cluster head acts as an agent for all members of the cluster, conducts local model training, and communicates with the server. Each cluster member, except the head, sends a distilled data of much smaller size to the cluster head. In this work, clients employ the KIP method to obtain distilled data, as shown in Section III-A. For the $h$-th heterogeneous cluster, when the cluster head aggregates the distilled data from the entire cluster, it obtains a dataset $\mathcal{D}_{h}$, whose distribution approximates the global distribution. We transform the problem into a learning process between the server and a number of cluster heads with approximately IID data.
@@ IV-B4 Model Training
As we mentioned above, after transforming our problem into distributed learning with IID data, we can apply the classical FL method, such as FedAvg[11], for model training. In the communication round $t$, the server broadcasts the global model parameters $\omega^{t}$ to all heads of the connected cluster. Subsequently, the cluster head $h$ updates its local parameters to $\omega_{h}^{t}\,\,=\,\,\omega^{t}$, and performs $E_{2}$ local updates in $\mathcal{D}_{h}$. At the $(k+1)$-th step of local update, the local model parameters of cluster head $h$ are updated as
@@
where $h=1,\cdots,H$, $k=0,\cdots,E_{2}-1$, $\xi_{h}^{t,k}$ denotes a randomly selected batch from $\mathcal{D}_{h}$, $\eta$ denotes the learning rate and $\nabla F_{h}\left(\cdot\right)$ denotes the gradient of the local loss function. Finally, the server aggregates the updated local models $\omega_{1}^{t,E_{2}},...,\omega_{H}^{t,E_{2}}$ from all the heads of the cluster to generate a new global model.
The general process of the proposed HFLDD is illustrated in Algorithm 2.
@@
19:     select cluster head $a_{h}$ randomly from $\mathcal{H}_{e,h}$
@@
V Performance Analysis
V-A Convergence Analysis
In HFLDD, the global training process is organized around a set of cluster heads ${h=1,\dots,H}$, each of which holds a hybrid dataset consisting of its own local data and the distilled data received from its cluster members. Each cluster head independently performs model training using this hybrid dataset, which approximates the distribution of the entire cluster. The convergence behavior of HFLDD can be analyzed by extending the standard convergence theory of FedAvg with non-IID data. HFLDD maintains the following foundational assumptions commonly adopted in FL:
Assumption 1 ($L$-smoothness) : For any model parameters $\upsilon$ and $\omega$,
$||\nabla f_{h}(\omega)-\nabla f_{h}(\upsilon)||^{2}\leq L||\omega-\upsilon||^{2},\forall h=1,\dots,H.$ (6)
Assumption 2 ($\mu$-strong convexity): For any model parameters $\upsilon$ and $\omega$,
$f_{h}(\upsilon)\geq f_{h}(\omega)+(\upsilon-\omega)^{T}\nabla f_{h}(\omega)+\frac{\mu}{2}||\upsilon-\omega||^{2},\forall h=1,\dots,H.$ (7)
Assumption 3 (Unbiased stochastic gradients): Let $\xi_{h}\sim\mathcal{D}_{h}$ be a mini-batch sample from the hybrid dataset of cluster $h$, the stochastic gradient satisfies:
$\mathbb{E}_{\xi_{h}\sim\mathcal{D}_{h}}\left[\nabla f_{h}(w,\xi_{h})\right]=\nabla f_{h}(w)$ (8)
$\mathbb{E}_{\xi_{h}\sim\mathcal{D}_{h}}\left[||\nabla f_{h}(w,\xi_{h})-\nabla f_{h}(w)||^{2}\right]\leq\sigma_{h}^{2}$ (9)
Assumption 4 (Gradient Boundedness): The expected squared norm of stochastic gradients is uniformly bounded:
$\mathbb{E}_{\xi_{h}\sim\mathcal{D}_{h}}\left[||\nabla f_{h}(w,\xi_{h})||^{2}\right]\leq G^{2},~{}\forall h=1,\dots,H.$ (10)
To quantify the degree of data heterogeneous across clusters and the approximation error of dataset distillation, we further make the following assumptions:
Assumption 5 (Bounded heterogeneity across clusters): Let $\bar{\mathcal{D}}_{h}$ denote the full dataset containing the raw data of all the clients in cluster $h$. Let the optimal solution of the global model be $\omega^{*}\triangleq\arg\min_{\omega}f(\omega)$, the minimum global loss be $f^{*}\triangleq f(\omega^{*})$, and the minimum local loss at cluster $h$ based on $\bar{\mathcal{D}}_{h}$ be $f_{h|\bar{\mathcal{D}}_{h}}^{*}$, then
$|f^{*}-\sum_{h=1}^{H}\mathrm{w}_{h}f_{h|\bar{\mathcal{D}}_{h}}^{*}|\leq\Gamma_{cluster},$ (11)
where $\mathrm{w}_{h}$ denotes the aggregation weight of local model at cluster $h$.
Assumption 6 (Bounded distillation error): The approximation error between the empirical loss on real dataset and the distilled data at each cluster head is bounded by:
$|f_{h|\bar{\mathcal{D}}_{h}}^{*}-f_{h|{\mathcal{D}}_{h}}^{*}|\leq\Gamma_{distill}.$ (12)
Based on the above six assumptions, we have the following theorem with regards to the convergence of HFLDD.
Theorem 1.
Under Assumptions 1–6, if the learning rate follows a diminishing schedule $\eta_{t}=\frac{2}{\mu(\tau+t)}$, with $\tau=\max\left\{8\frac{L}{\mu},\ E_{2}\right\}-1$, and all cluster heads participate in each communication round, then after $t$ rounds of communication, the global model $\omega_{T}$ satisfies:
$\mathbb{E}[f(\omega^{t})]-f^{*}$ $\leq\frac{L}{t+\tau}\left(\frac{2Q}{\mu^{2}}+\frac{\tau+1}{2}||\omega^{0}-\omega^{*}||^{2}\right),$ (13)
where
$Q=\sum_{h=1}^{H}\mathrm{w}_{h}^{2}\sigma_{h}^{2}+6L\left(\Gamma_{cluster}+\Gamma_{distill}\right)+8(E_{2}-1)^{2}G^{2}.$ (14)
Proof.
In HFLDD, only the cluster heads participant in global model training. It can be regarded as standard FedAvg on a cluster-level hybrid dataset $\mathcal{D}_{h}$. Based on the convergence analysis of FedAvg under non-IID conditions, as shown in [12], the convergence bound of HFLDD satisfies:
$\mathbb{E}[f(\omega^{t})]-f^{*}$ $\leq\frac{L}{t+\tau}\left(\frac{2B}{\mu^{2}}+\frac{\tau+1}{2}||\omega^{0}-\omega^{*}||^{2}\right),$ (15)
where
$B=\sum_{h=1}^{H}\mathrm{w}_{h}^{2}\sigma_{h}^{2}+6L\Gamma+8(E_{2}-1)^{2}G^{2}$ (16)
The parameter $\Gamma=f^{*}-\sum_{h=1}^{H}\mathrm{w}_{h}f_{h|{\mathcal{D}}_{h}}^{*}$ quantifies the degree of non-IID across datasets ${\mathcal{D}}_{h},h=1,\cdots,H$. Under the assumptions of bounded heterogeneity and distillation approximation error, we have
$\Gamma$ $=\underbrace{f^{*}-\sum_{h=1}^{H}\mathrm{w}_{h}f_{h|{\bar{\mathcal{D}}_{h}}}^{*}}_{\text{cluster heterogeneity}}+\underbrace{\sum_{h=1}^{H}\mathrm{w}_{h}\left(f_{h|{\bar{\mathcal{D}}_{h}}}^{*}-f_{h|{{\mathcal{D}}_{h}}}^{*}\right)}_{\text{distillation error}}$
$\leq\left|f^{*}-\sum_{h=1}^{H}\mathrm{w}_{h}f_{h|{\bar{\mathcal{D}}_{h}}}^{*}\right|+\sum_{h=1}^{H}\mathrm{w}_{h}\left|f_{h|{\bar{\mathcal{D}}_{h}}}^{*}-f_{h|{{\mathcal{D}}_{h}}}^{*}\right|$
$\leq\Gamma_{cluster}+\Gamma_{distill}$ (17)
Combining Eq.(16) and Eq.(V-A), we can easily derive that $B\leq Q$. Straightforwardly, Eq. (13) holds. This concludes the proof.
Theorem 1 demonstrates that under Assumption 1-6, when using a diminishing leaning rate $\eta_{t}=\mathcal{O}(1/t)$, the expected global loss converges to the optimal value at a rate of $\mathcal{O}(1/t)$. Moreover, the convergence bound is closely related to the heterogeneity of the cluster and the distillation error. The higher the heterogeneity of the cluster or the distillation error, the larger the convergence bound.
V-B Communication Cost
Now, we analyze the total communication cost of HFLDD. Considering that the initialization of the server model can be sent as a random seed [29], we can ignore the communication cost of sending random seeds. Thus, we only consider the resources consumed for uploading soft labels to the server at the stage of label knowledge collection, transmitting the distilled data within each cluster at the stage of IID dataset generation, and communication between the server and cluster heads at the stage of model training.
Let $a_{h}$ denote the head of the heterogeneous cluster $h$, and $\left|\tilde{\mathcal{D}_{i}}\right|$ denote the size of the distilled dataset on the client $i$. Suppose that transmitting each parameter incurs a communication overhead of $B_{1}$ (usually 32 bits), and transmitting each distilled data incurs a communication overhead of $B_{2}$ (1 × 8 bits or 3 × 8 bits for each grayscale or RGB pixel). The bit length $B_{2}$ reflects the transmission cost of each image sample, and no additional compression beyond the inherent reduction via dataset distillation is applied. For $N$ clients, the communication cost for uploading the soft label is $N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}$. The total communication cost to transmit the distilled data within each cluster is $\sum_{h=1}^{H}{\sum_{\begin{subarray}{c}i\in\mathcal{H}_{e,h}\\ i\neq a_{h}\end{subarray}}{\left|\tilde{\mathcal{D}_{i}}\right|}}\cdot B_{2}$. In the model training stage, the total cost of model communication between the server and the heads of the cluster in $T$ rounds is $H\cdot|\omega|\cdot(2T-1)\cdot B_{1}$. Consequently, the total communication cost of HFLDD is
$Cost_{hfldd}$ $=N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}$
$~{}~{}~{}+H\cdot|\omega|\cdot(2T-1)\cdot B_{1}.$ (18)
V-C Computational Complexity
The computational complexities of HFLDD at the server, the cluster members and the cluster heads are separately analyzed.
V-C1 Server
In HFLDD, the server is responsible for client clustering and global model aggregation. In the process of client clustering, the server needs to compute the similarity matrix $M$, which involves the computation of KL Divergence between any pair of clients’ soft labels. As shown in Eq. 4, for each pair of clients $i$ and $j$, the complexity of computing $M_{ij}$ over the global dataset $D_{g}$ is $\mathcal{O}(|D_{g}|\cdot N_{c})$. In total, the complexity of constructing the similarity matrix $M$ is $\mathcal{O}(N^{2}\cdot|D_{g}|\cdot N_{c})$. Based on the similarity matrix $M$, the server conducts homogeneous client clustering using K-Means algorithm, which incurs a complexity of $\mathcal{O}(I\cdot K\cdot N^{2})$, where $I$ is the number of iterations in K-means. In the process of global model aggregation, the server aggregates models from all $H$ cluster heads over $T$ communication rounds. Each aggregation involves weighted averaging $H$ local models with size being $|\omega|$, resulting in a total computational cost of $\mathcal{O}(T\cdot H\cdot|\omega|)$.
V-C2 Cluster Member
Each cluster member performs pre-training and soft label generation at the stage of label knowledge collection and dataset distillation at the stage of IID dataset generation. In the process of local pre-training, each client performs $E_{1}$ steps of model updates on its local dataset using mini-batch SGD. Let $b_{1}$ denote the mini-batch size used in each step. The computational complexity of local pre-training per client is $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|)$. After pre-training, each client generates a soft label tensor by performing inference over the public dataset $\mathcal{D}_{g}$, which incurs an additional computational cost of $\mathcal{O}(|\mathcal{D}_{g}|\cdot|\omega|)$. Therefore, the computation cost per client at the label knowledge collection stage is $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|+|\mathcal{D}_{g}|\cdot|\omega|)$. In the process of dataset distillation, each cluster member synthesizes distilled data via the KIP method, which involves repeated optimization over $R$ iterations. In each iteration, for cluster member $i$, $i\in\mathcal{H}_{e,h}$ and $i\notin\mathcal{A}$, the most computationally intensive step is the inversion of the regularized kernel matrix in kernel ridge regression, which incurs a cost of $\mathcal{O}(|\tilde{\mathcal{D}_{i}}|^{3})$. Therefore, the complexity of dataset distillation at cluster member $i$ is $\mathcal{O}(R\cdot|\tilde{\mathcal{D}_{i}}|^{3}).$
V-C3 Cluster Head
In HFLDD, at the stage of label knowledge collection, each cluster head follows the same pre-training and soft label generation steps as cluster members, which incurs a computational complexity of $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|+|\mathcal{D}_{g}|\cdot|\omega|)$. At the stage of global model training, each cluster head additionally conducts local training on its hybrid dataset. In each communication round, the cluster head at cluster $h$ performs $E_{2}$ steps of mini-batch SGD on its dataset $\mathcal{D}_{h}$. Let $b_{2}$ denote the batch size. The computational complexity of local training per round on a single cluster head is $\mathcal{O}(E_{2}\cdot b_{2}\cdot|\omega|)$. For all $T$ rounds, the total complexity is $\mathcal{O}(T\cdot E_{2}\cdot b_{2}\cdot|\omega|)$.
VI Experimental Evaluation